find the intergral solutions(x,y) of the equation 
2xy – 4x2 + 12x – 5y = 11

Grade:12

1 Answers

vsujji
8 years ago
4x^2-12x+11=2xy-5y
 [2x]^2-2[2x]3+9-9+11=2xy-5y
 [2x-3]^2+2=2xy-5y
 [2x-3]^2+2=y[2x-5]
    so by trial and error method we get [x,y]=[3,11] ;[4,9]